Particular integral for cos^3(x÷2)?Could you please tell me sir?

Grade:12th pass

1 Answers

Anish Singhal
5 years ago
Split cos into two parts one with exponent 2 and another with exponent 1 and then write cos^2(x/2)=1-sin^2(x/2), then use substitution sin(x/20=t and proceed.